A comet is traveling through space with speed 3.73 ✕ 104 m/s when it encounters an asteroid that was at rest. The comet and the asteroid stick together, becoming a single object with a single velocity. If the mass of the comet is 1.21 ✕ 1014 kg and the mass of the asteroid is 6.30 ✕ 1020 kg, what is the final velocity of their combination? (Assume the comet's initial direction is positive. Indicate the direction with the sign of your answer.)

Vf=?
Answer #1

using conservation of momentum

m1 v1 + m2 v2 = (m1 + m2) v

1.21* 10^14* 3.73*10^4 + 0 = ( 1.21 *10^14 + 6.3 * 10^20)v

v = 7.164*10^-3 m/s

=========

Comment in case any doubt, will reply for sure.. Goodluck